OpenStudy (anonymous):

okay, so this is probably a stupid question, but im just not sure.. i think its A, but am i right? what confuses me is the other choices are very different and i just dont know. What is the prime factorization of 72? A. 8*9 B. 2^3*3^2 C. 2^3*9 D. 2^4*3^3

OpenStudy (bibby):

Make a tree and break down the numbers into their respective factors like so: 

OpenStudy (anonymous):

okay so im right?

OpenStudy (bibby):

no, that's why I drew the lines below the 8 and 9 they can both be factored further

OpenStudy (anonymous):

ooohhhhhhhhh so its D. 8 = 2^4 and 9 =3^3

OpenStudy (anonymous):



OpenStudy (anonymous):

oh

OpenStudy (bibby):

2^4 is 16 :p

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

How many 2s are in my drawing and how many 3s?

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhh

OpenStudy (bibby):

You're supposed to factor the numbers until you have prime numbers. like 2,3, 5, 7 etc. Then you rewrite the number as a product. 2*2*2*3*3*3*3

OpenStudy (anonymous):

3*3*2*2*2 3*3=3^2 2*2*2=2^3 Notice? So yes exactly what @bibby said.

OpenStudy (anonymous):

okay, i get it. haha thank you, both.